### v2

# Scaling Laws for Function Diversity and Specialization Across Socioeconomic and Biological Complex Systems

Abstract
Function diversity, or the range of tasks that individuals perform, is essential for productive organizations. In the absence of overarching principles, the characteristics of function diversity are seemingly unique to each domain. Here, we introduce an empirical framework and a mathematical model for the diversification of functions in a wide range of systems, such as bacteria, federal agencies, universities, corporations, and cities. Our findings reveal that the number of functions within these entities grows sublinearly with system size, with exponents ranging from 0.35 to 0.57, confirming Heaps’ Law. In contrast, cities exhibit logarithmic growth in the occupation types. We generalize the Yule-Simon model to quantify a wide range of these empirical observations by introducing two new key attributes: a diversification parameter that characterizes the tendency for more populated functions to inhibit new function creation, and a specialization parameter that describes how a function’s attractiveness depends on its abundance. These parameters allow us to position diverse systems, from microorganisms to metropolitan areas, within a two-dimensional abstract space. This mapping suggests underlying commonalities and differences in the foundational mechanisms that drive the growth of these systems.
1 Introduction
Fundamental to complex adaptive systems—encompassing biological organisms, human organizations, and urban areas—is the range of functions (diversity) and their relative abundance (specialization). Diversity, or the composition of a system’s sub-components, indicates a broad range of abilities and capabilities for organisms or organizations to tackle complex challenges, hedge against risks, and thus adapt to complex environments [1, 2, 3, 4, 5, 6]. As these systems mature, develop, and expand, their function diversity grows, reflecting an increase in adaptability and resilience [7, 1, 2, 8, 3]. Remarkably consistent empirical regularities are observed in both biological and socio-economic systems, despite their vastly different environmental conditions and evolutionary trajectories. Examples include the species-area curve—the empirical regularity for how the number of plant and animal species increases with land area [2]—and the scaling of microbial richness with community size [9, Fig. 1]. A similar pattern is observed in urban areas—the range of occupations and business types is strongly associated with population [4, 5, 10].
While diversity provides a wide foundation of capabilities, specialization refines these capabilities for targeted functions, concentrating on a select few functions or skills to increase performance and competitive edge [11, 12]. For example, certain flowers adapt by developing nectars tailored to their specific pollinators, universities may concentrate on certain research areas like quantum computing, and cities may become hubs for certain sectors such as financial services. Specialization is manifested in the abundance distribution of functions—whether components or resources are concentrated in a few functions or evenly spread across many. Empirically, specialization has been quantified using indices like the Gini coefficient, location quotient, and entropy measures, all of which capture information in the abundance distribution [13, 14, 15, 12]. Specialization within the diversification process provides insights into how systems efficiently allocate their resources, stand out in competitive environments, and add value to the broader system’s complexity and resistance [16].
Indeed, different systems pursue distinct strategic choices, objectives, and impacts under different specific scopes and subject matters. Accordingly, studies of diversity and specialization have predominantly been siloed within isolated disciplinary boundaries, thus limiting the possibility of a comprehensive understanding. Research efforts have often focused on specifics, ranging from the effects of team diversity [1, 17] to the ramifications of organizational diversity [18, 19, 20], labor specialization [21, 22], cellular heterogeneity [23, 24], or economic diversity and complexity [25, 26]. While these studies are invaluable within their respective fields, findings are often limited to a particular organizational type or the impact of single variables. This segmented approach runs the risk of missing out on identifying overarching principles and theories that underlie the processes of diversification and specialization across a variety of systems.
Thus there is a need for a comprehensive framework to study diversification and specialization across complex systems. Evidence supports broad regularities across a wide range of systems. Previous studies have identified universal patterns of abundance fluctuations across microbial communities, tropical forests, and urban populations [27]. Additionally, Heaps’ Law, which describes regularities in the scaling of novel elements, has been observed in various human-generated artifacts, including books and musical compositions, and biological components [28, 29].
The existing literature in both economics [30, 31] and biology [32, 33, 34, 35] often presents diversification and specialization as opposing poles of a linear spectrum. While this perspective is valid, especially in the context of finite resource allocation, it does not universally hold. For instance, research has demonstrated that introducing more minor entities into a market can, paradoxically, reinforce the dominance of the major players [36]. This indicates that diversification and specialization may be more accurately understood as a two-dimensional problem.
In this study, we investigate function diversity and specialization in a broad range of socio-economic and biological systems, including bacteria, companies, government agencies, universities, and cities. We explore the interconnected nature of diversification and specialization across a wide range of systems by asking: Do various biological and social systems display common patterns of function diversity and abundance, and to what extent? What general governing principles could determine the function diversity and specialization of complex adaptive systems? We first measure the range (diversity) and concentration of functions in these systems (abundance distributions) within each system. We identify functional forms for each attribute to compare disparate systems. We then propose a general model for the growth of function diversity and abundance, explaining both the regularities and differences observed in the empirical data.
Our cross-system perspective recognizes that biological and social systems exhibit common patterns of complexity, adaptation, and evolution. Certain principles, like network dynamics [37] and feedback loops [38], are common across complex systems. Exploring common governing principles across systems may reveal underlying universal patterns that govern complex systems in general. This view is supported by several prior studies that have found commonalities in related domains, such as in the distribution of business types in cities [4, 5], and the abundance fluctuations across biological, ecological, and urban systems [27].
2 Results
2.1 Empirical Results
Figure 1: The number of distinct functions versus system size (top row) and rank-frequency distribution of function abundance (bottom row) in several complex adaptive systems. (A,D) Bacterial and archaea cells, (B,E) US federal agencies, and (C,F) US metropolitan statistical areas (MSA). The dependence of the number of functions on size measure $N$ in (A) and (B) can be approximated by the power law, $D\sim N^{\beta}$. Also shown is the best-fit scaling exponent, $\beta$, and its $95\%$ confidence interval in the square brackets. (C) In contrast, logarithmic scaling occurs for MSA’s, $D\sim\log N$.
Table 1: Summary of scaling statistics for function diversity in various biological and social complex adaptive systems
Size measure Function diversity measure $\beta$ 95% CI R-sq
Bacteria Total expressed protein Distinct expressed proteins 0.35 [0.31, 0.39] 0.86
Norwegian companies Number of employees Distinct occupations 0.46 [0.45, 0.46] 0.88
US federal agencies Number of employees Distinct occupations 0.50 [0.46, 0.53] 0.87
US universities, associate level Number of faculty Distinct academic programs 0.52 [0.50, 0.55] 0.61
US universities, bachelor level & above Number of faculty Distinct academic programs 0.57 [0.55, 0.59] 0.75
Cities Working population Distinct occupations Logarithm scaling 0.94
Our study analyzes empirical data from microorganisms to firms to metropolitan areas, including 47 bacteria and archaea cell observations [39], 125 US federal government agencies, 3,191 Norwegian companies, 2,397 US universities, and 422 US metropolitan areas. For each system, we determine the diversity, the number of distinct functions $D(N)$ and the distribution of these functions’ abundance $f_{i}$ as a function of system size $N$.
We define system size as the total number of individuals within a system. In social systems, this is simply the number of people in the system, with functions represented by occupations, which are categorized using standardized classifications (see SI for details on data and methods). In bacteria and archaea, key functional components responsible for metabolism and cellular processes are proteins, analogous to workers in an organization. Accordingly, we measure system size in these biological systems as the total number of expressed proteins, while function is quantified as the number of distinct expressed proteins. Figure 1 illustrates how the range of distinct functions, $D(N)$, scales with the system size $N$. This scaling framework allows for direct comparisons across diverse systems, as demonstrated in previous studies utilizing scaling analysis [40, 41, 42]. The data for bacteria and archaea are visualized in panel A, federal agencies in panel B, and metropolitan areas in panel C. Analysis of additional datasets, including Norwegian companies and universities, are summarized in Table 1.
For most systems, this scaling behavior is well-captured by a power-law relation $D=D_{0}N^{\beta}$, where $\beta$ is the scaling exponent, and $D_{0}$ is a constant. For bacteria and archaea, companies, federal agencies, universities of associate level, bachelor level and above, the scaling exponents are sublinear, ranging between $0.35$ to $0.57$. Notably, such scaling behavior resembles Heaps’ Law, which is observed in the growth of unique types across diverse human artifacts and biological components [28, 29]. Across bacteria and archaea, the consistency of this pattern is striking, given that each of the proteomes comes from organisms that have followed unique evolutionary trajectories, have different metabolisms, and occupy unique ecological niches. Comparative analyses of systems ranging from books and musical compositions to genomes, using appropriate diversity metrics, have consistently identified scaling exponents between 0.35 and 0.60, showcasing a consistent scaling pattern across both cultural artifacts and biological systems [43, 44, 45, 46, 47].
Among the systems analyzed, a notable exception is metropolitan areas. Unlike other datasets, the scaling of function diversity in metropolitan areas does not follow a power-law pattern but is better captured by a logarithmic relationship, $D=b\log N+c$ , as shown in Figure 1(C). This observation suggests that different mechanisms govern the growth of function diversity in cities compared to other systems. However, the sub-linear growth of all of these systems implies that the rate of increase of their function diversity systematically slows down as the system grows in size. For example, this pattern indicates that function diversity grows more rapidly with population increase in smaller cities but decelerates as cities become larger [4].
This logarithm scaling behavior carries two important implications for urban areas. First, if a city’s population grows exponentially with time ($t$), i.e., $N(t)=N_{0}e^{\delta t}$, where $N_{0}$ is the population at time $t=0$, and $\delta$ is the growth rate, then function diversity is predicted to grow linearly in time, with $D=\delta bt+D_{0}$, where $b$ is the parameter fitted from $D=b\log(N)+c$, and $D_{0}=b\log(N_{0})+c$ is a constant determined by the initial condition of the system at $t=0$. Then the growth rate of function diversity ($dD/dt$) is given by the product, $\delta\,b$, which is proportional to the population growth rate. For example, the growth rate $\delta$ for US urban areas on average is $0.71\%$ in 2019 [48] while the fitted parameter $b$ from data is $128$, which together predict that the growth rate for function diversity ($dD/dt$) is $\delta\,b=0.9$ occupations per year. Second, since most socio-economic metrics in urban areas, such as GDP, wages, and patent production, follow an approximate superlinear power-law relationship with population size [41], our findings suggest that function diversity serves as a key driver of recombinant growth [49, 50]. As functional diversity expands, the combinatorial possibilities for innovation and productivity multiply, contributing to exponential increases in economic output. This underscores the central role of diversity in sustaining and accelerating economic performance in urban systems.
Lastly, we analyze the distribution of function abundance within each organization. Figure 1(D), (E), and (F) illustrate the variation in the relative abundance of functions as a function of rank for bacteria and archaea, federal agencies, and metropolitan areas, respectively. Grey dots represent all data points, while select organizations are highlighted in color for clarity.
Despite variation in the most prevalent functions across organizations—such as “nurse” and “medical officer” dominating in Veterans Affairs, while the National Science Foundation’s top occupations include “miscellaneous administration and program” and “management and program analysis”—the rank-frequency distributions exhibit a consistent concave shape across all systems. This pattern suggests an exponential decline in function abundance with increasing rank. Notably, the distribution in metropolitan areas displays a striking level of uniformity compared to the more heterogeneous patterns observed in federal agencies and cellular systems. This consistency in urban areas aligns with previous research highlighting similar diversity patterns in business types across cities [4, 10].
Our analysis reveals both commonalities and differences across systems. The commonality is manifested in the fact that within a given system (bacteria, companies, etc.), the range of function diversity generally grows as a sublinear power law with system size. This commonality suggests that while varying in goals, environment, and history, different organisms may share a similar underlying dynamic in how their function range is generated and grows with size. The differences across systems, however, are also important and manifested in two ways. First, the function range in cities scales differently from the other systems, suggesting the mechanisms that lead to diversity in cities may be fundamentally different from the other entities. Second, among the systems that exhibit power law scaling, there are variations in the mechanisms that give rise to the different exponents. Our analyses are the first to attempt to quantitative assessment for comparison across disparate systems.
These findings present both challenges and opportunities for a comprehensive theory that can account for both the observed commonalities and differences. In particular, while existing models have typically predicted either power-law or logarithmic scaling of function diversity (e.g., [51, 4]), none have successfully integrated both patterns. In the next section, we propose a model that is capable of explaining the full spectrum of empirical observations related to function diversity and specialization. By identifying and incorporating key model parameters, our approach aims to shed light on the diverse mechanisms at play, offering a unified understanding of function diversity and specialization across systems.
2.2 Mathematical Model
Our model aims to explain the growth dynamics of a system as new individuals join. Within this framework, a new member that joins the system can lead to one of two scenarios: (i) the new member creates a new function (diversification) or (ii) the newcomer assimilates into a pre-existing function (specialization). For example, the individuals entering the system are: newly synthesized proteins in bacteria; newly hired staff in federal agencies; and new professionals entering the job market in cities. These dynamics are aligned with the foundational principles of the Yule-Simon model [52, 53]. We mathematically represent these dynamics by postulating the following growth rule for $n_{k}$, the number of functions with $k$ individuals:
$\frac{\partial n_{k}}{\partial N}=p\delta_{k,1}+(1-p)\left(q_{k-1}n_{k-1}-q_{k}n_{k}\right)\;.$ (1)
Here $p$ is the probability of a new function being created when a new individual joins the system, $q_{k}$ denotes the probability that the new individual joins a function that currently contains $k$ individuals, and $\delta_{k,1}$ is the Dirac delta function. The first term of the equation accounts for the creation of a new function, in which the number of functions with $k=1$ individuals increases by 1. The second term accounts for an individual that joins an existing function. If the individual joins a function with $k-1$ existing individuals, $n_{k}$ increases by 1. If the individual joins a function with $k$ individuals, $n_{k}$ decreases by 1 (as that function now has $k+1$ individuals). In other scenarios, $n_{k}$ remains unchanged. Figure 2 visually summarizes this dynamic process.
Figure 2: A conceptual diagram illustrating the structure of the mathematical model. Functions are characterized by the number of individuals in them. When a new function is created, it adds to functions with 1 individual (with probability $p$ in each time step). Functions with $k$ individuals move to the next category as an individual joins this function (with probability $q_{k}$ in each time step).
Unlike the classic Yule-Simon model, which assumes a constant rate $p$ of new function creation, our approach introduces a dynamic framework. As a system adapts to a changing environment, new functions may become necessary. However, as the number of existing functions increases, it becomes more likely that the needed functionality is already present. Consequently, the rate of new function creation decreases. When the system contains all possible functions, this rate should approach zero. A simple functional form that would capture the decreasing tendency of $p$ is $p=p_{0}/D$. The constant $p_{0}$ represents effects from environmental factors that are assumed to be the same across systems of the same type.
More generally, the rate of new function creation may not scale inversely with diversity alone, but instead depend on the system’s internal structure, as captured by its rank-frequency distribution. A flexible parameterization that accounts for this is,
$p=\frac{p_{0}}{\sum_{i=1}^{D}k_{i}^{\theta}}\;.$ (2)
where $k_{i}$ denotes the number of individuals within function $i$. The parameter $\theta$ modulates the influence of function abundance on the creation of new functions. When $\theta=0$, we recover the earlier linear dependency, $p=p_{0}/D$. For $\theta>0$, functions with more individuals more strongly suppress the addition of new functions. For $\theta<0$, the opposite holds. This generalized form allows for a wide range of dynamics to be modeled at the level of individual systems.
When a new individual joins an existing function, we consider the probability of the individual joining each function to be affected by the abundance of the existing functions. Having more individuals in a function can increase the likelihood of a new individual joining this function. In some scenarios, it may have the opposite effect because the existing abundance already adequately meets the system’s demands. We can formulate this tradeoff as a nonlinear preferential attachment process—that the probability of an individual joining an existing function with $k$ individuals, $q_{k}$, is
$q_{k}=\frac{k^{\gamma}}{\sum_{i=1}^{D}k_{i}^{\gamma}}\;,$ (3)
where $\gamma$ is a specialization parameter that indicates the extent to which large functions attract newcomers. This equation is similar to previous models of preferential attachment in physics [54, 55, 37] and circular causation [56], positive feedback [57] or agglomeration [58] in economics. Our model introduced $\gamma$ as a parameter, ranging from 0 to 1, to control the strength of this positive feedback. For example, on the one hand, sublinear preferential attachment [37], that is, $\gamma<1$, denotes diminishing returns in the effect of function abundance on joining an existing function. On the other hand, the cases $\gamma=1$ indicate linear preferential attachment, and $\gamma>1$ corresponds to superlinear preferential attachment, which is extremely sensitive to the existing size of the function .
In summary, our model contains two key parameters, $\theta$ and $\gamma$, that characterize the effect of diversification and specialization, respectively. While $\theta$ primarily determines the expansion of the function range, $\gamma$ primarily shapes the accumulation of individuals in functions.
2.3 Model predictions and categorization of systems by diversification and specialization parameters
We simulate a system’s growth based on Eq. 1 and compare the model’s predictions with empirical data. We then estimate the parameters $\gamma$ and $\theta$ for each system by comparing the function range and abundance generated by simulation with those from data on bacteria, federal agencies, and cities. This estimation is achieved by minimizing the Euclidean distance between normalized rank-frequency distributions in logarithmic-space using the adaptive differential evolution algorithm for optimization [59] (see SI for methods).
Figure 3 (a-c) shows the model’s predicted rank-frequency distributions are well aligned with three selected systems: Bartonella henselae (bacterial cell), the Army (federal agency), and Warren-Troy-Farmington Hills (urban area in Michigan). See SI for more examples of rank-frequency distribution predictions. The bottom row of Fig. 3 show their respective diversity relationship with those predicted by the model. Both results demonstrate good agreement across all three types of systems for both the rank-frequency distributions and the diversity scaling.
Figure 4 summarizes our simulation results, mapping over a hundred cases onto the parameter space of $\theta$ and $\gamma$. Systems cluster by type, indicating distinct within-class dynamics. At the same time, patterns emerge across systems along both $\theta$ and $\gamma$, revealing shared mechanisms that operate both within and across system classes. For instance, bacteria, federal agencies, and cities exhibit parallel trends in the specialization parameter $\gamma$, which captures how the abundance of existing functions shapes subsequent growth. All three systems consistently show positive $\gamma$ values, but predominantly sublinear (below 1), indicating a general tendency toward diminishing returns in specialization—growth is less likely to concentrate solely on already dominant functions. Nevertheless, the variation in $\gamma$ is wider for federal agencies and bacteria than for cities, suggesting that urban systems may exhibit a more uniform—and possibly universal—specialization pattern. On the other hand, $\theta$ has a broader range for different class of systems than $\gamma$.
Figure 3: Summary of the model calibration results for the normalized rank-frequency distributions of (a) bacterial cells, (b) federal agencies and (c) cities, each for a given case of that system. The legends show the calibrated values of $\gamma$ and $\theta$, as well as the particular system in question. Each system was simulated using an initial condition that arose from one of the lesser size organizations in that data set (see SI for further details). Panels (d), (e), and (f) show the respective diversity plots from simulations given the mean values of $\theta$ across all calibrations for that system, starting from the initial condition size used for calibration.
While the specialization parameter, $\gamma$, exhibits common patterns across systems in terms of its mean value, the diversification parameter, $\theta$, differs significantly across systems. It is notably higher for cities than for federal agencies and cells. A higher $\theta$ value suggests the existence of a highly abundant function suppresses new function creation. This is in line with the observation that large cities tend to become hubs for certain industries [60], such as San Francisco Bay Area for information technology and the Greater Boston area for biotech. A value of $\theta$ around zero implies that an increased range of existing functions, regardless of the size of the functions, contributes towards suppression of new functions. The differences in $\theta$ parameters can help explain the differences in scaling behavior observed. When $\theta=1$, given that $\sum_{i=1}^{D}k_{i}=N$, we can simplify Eq. 2 to derive $p=p_{0}/N$. Since $p=dD/dN$, we arrive at the dynamical equation $dD/dN=aD_{\text{max}}/N$. The solution of this equation takes the form of $D\sim\log(N)$, which recovers the logarithmic scaling. Similarly, in the case of $\theta=0$, we have $\sum_{i=1}^{D}k_{i}^{0}=D$. This leads to $dD/dN\sim 1/D$, which implies the power-law scaling $D\sim N^{1/2}$. For bacteria, Figure 1(A) shows that diversification happens faster than in cities, but slower than in federal agencies, and correspondingly the mean calibrated value of $\theta$ is inbetween 0 and 1.
Figure 4: The parameter space of specialization ($\gamma$) and diversification ($\theta$) estimated for each system across different classes of systems. The vertical dashed line shows $\theta=1$, the horizontal dashed line shows $\gamma=1$, and the diagonal dashed line shows $\gamma=\theta$. We selected the federal agencies with $N>1000$ (40 instances), all of the bacterial data we had access to (46 species), and the largest twenty cities.
Each class of systems occupies distinct regions of parameter space. First, the values of $\theta$ found for cities imply that diversity growth is most impeded when large functions are abundant. Second, for federal agencies, smaller $\theta$ values imply that functions more equitably impede diversity growth. Finally, for bacterial cells, an intermediate $\theta$ value implies an in-between level of diversity growth still impeded by large functions. Cities express the largest value of $\theta$ possible (approximately 1) that simultaneously allows for the maximization of concentration inzq few functions and universality of the rank-frequency distributions without leading to gelation behavior that occurs for $\gamma>1$ [37]. In this sense, gelation provides a natural bound to function growth since it does not allow for the robustness or diversity necessary for organizations to survive complex environments.
3 Discussion
We identify common patterns in the diversity and abundance of functions across a broad spectrum of biological and social complex systems. In systems such as bacteria, federal agencies, companies, and universities, function diversity can be approximated by a power-law relationship with size, characterized by sublinear exponents ranging from 0.35 to 0.57, depending on the system. This is consistent with Heaps’ Law. However, the function diversity of cities follows a different pattern, scaling logarithmically with size.
We propose a general model for the growth of function diversity and abundance that accounts for both power-law and logarithmic scaling behaviors. The model is characterized by two key parameters: a diversification parameter, which governs the expansion of function range, and a specialization parameter, which grows function concentration. We estimate both parameters from the data for bacterial cells, federal agencies, and cities. Notably, we observe consistent sublinear preferential across all three systems, with generally small variations in the specialization parameter. However, the diversification parameter varies significantly across systems, suggesting different drivers for function range expansion. In our model, the probability of new function addition is, by definition, a decreasing function of the size of the organism. This choice is motivated by research indicating that not only is this the case in biological systems [61, 62], but more recently as a principle of ecology in the decreasing relationship between speciation rates and animal size (see [63, Figure 3]).
Our framework offers a complementary alternative to several influential theories in biology and ecology that have sought to explain the origins and patterns of diversity. Classical models such as the Island Biogeography Theory [64], the Unified Neutral Theory of Biodiversity and Biogeography (UNTBB) [65, 66], and Maximum Entropy [67] approaches have all been pivotal in modeling biodiversity through the lens of ecological equilibria, spatial constraints, or information-theoretic principles. In particular, the UNTBB has offered a controversial but widely used null model, helping researchers benchmark observed diversity patterns across ecosystems and even genomes. Unlike these models, which are typically grounded in ecological or evolutionary narratives specific to biological systems, our approach derives from a generalized process-based model that incorporates both diversification and specialization dynamics, and it applies equally well to socio-economic systems such as cities and federal agencies. This broader scope enables cross-domain comparisons and identifies shared underlying mechanisms, while also capturing key differences—such as the logarithmic versus power-law scaling of diversity—in a unified parameter space.
There are several limitations to the proposed model. First, the choice of a modified Yule-Simon type process to model diversity growth is likely not unique. Compared to previous models [51], our model contains an extra parameter, which effectively allows modulation between the probability of adding new functions compared to the strength of preferential attachment. This is necessary since it does not seem possible to unify the dynamics of diversity for federal agencies, cells, and cities using a model with fewer than two parameters. The classic Yule-Simon model, which assumes a constant rate of new function addition, leads to a linear scaling of diversity with system size which is not observed in the systems we consider (which are either sublinear power-law or logarithmic). Frameworks under other narratives, such as urn models or sample-space reducing mechanisms, could potentially also be adapted with an extra parameter for the same purpose. While our model considers the addition of individuals, individuals also leave organizations and cities, just as genes can naturally be removed from genomes through evolution. It could be useful for future research to consider the removal of individuals in complex systems and how it would affect diversification and specialization. Additionally, our model assumes a direct proportionality between system size and age, though this may not always hold. However, the relationship is generally expected to be monotonically increasing on average, as described by Cope’s rule [68]. Our proposed mechanism is an effective mechanism that can describe the emergence of structures from a unifying, albeit simplified, framework. This provides a way to produce the diverse observed empirical pattern across systems in a minimal way, to observe whether coordination problems across biology and society utilize similar solutions.
Cities differ significantly from bacterial cells and federal agencies across various empirical measures. Unlike the power-law scaling observed in bacteria and federal agencies, cities exhibit logarithmic scaling in diversity. Additionally, cities display greater universality in abundance distributions. This divergence may be attributed to the fact that cities are structurally distinct from the other two systems. Federal agencies and bacterial cells function as integrated units with well-defined boundaries, akin to organisms, and the goal for adapting to its environment is on the level of each organism. In contrast, cities resemble ecosystems, lacking a singular objective, with diverse entities operating together, each primarily based on self-interest rather than collective adaptation. This fundamental difference in structure and goal underpins the unique scaling behaviors observed in cities, distinguishing them from the more organism-like dynamics of bacterial cells and federal agencies.
Our paper makes several contributions to the literature. First, in contrast to the prevalent trend of studying diversification and specialization within specific system types, our work presents an integrated analysis across a diverse array of systems, encompassing both biological (bacteria) and social systems (companies, government agencies, universities, cities). This approach identifies commonalities across many biological and social systems, while also identifying notable exceptions. It provides testable hypotheses to explore patterns in other systems. Second, while previous models explained power-law or logarithmic scaling relationships using distinct formulations and assumptions, we develop a general model that explains both types of scaling within one unifying framework. Third, by estimating parameters in the framework of this mathematical model, we can quantify common characteristics across complex systems and evaluate how systems differ from each other.
Acknowledgements
The authors would like to acknowledge the support of the National Science Foundation Grant Award Number EF–2133863. We would like to thank the UNM Center for Advanced Research Computing, supported in part by the National Science Foundation, for providing the research computing resources used in this work, in particular, Matthew Fricke. H.Y. acknowledges the Emergent Political Economies grant from the Omidyar Network, the NRF Global Humanities and Social Sciences Convergence Research Program (2024S1A5C3A02042671) and the support from the Institute of Management Research at Seoul National University. J.H. acknowledges the support of a Lou Schuyler grant from the Santa Fe Institute.
Code availability
The code necessary to replicate the results of the calibrations, and simulate the model, is available at https://github.com/jamesholehouse/Calibrations-for-Function-Diversity.